Please ignore the circled choices. I am double checking my work.
Find the area ynder the normal curve to the right of 2-67 a) 7486 4214 813 1056 e) .7357 α) .7486 b) .2514 с).8413 0.1056 (9.7357 proportion of soores less than 18. Find the area under the noemal curve between -1.25 and-1. 19. Ir 10 scores are nomally distibuted with a mean o 20. Same as above except find the proportion of sooees greater 21. Consider a university with a population mean GPA of 2.95, a standard deviation of 100 and a standard deviation of 15, find the а).3173 b) .6827 @.2119 d).0359 e).I587 α) .3173 b),6827 u2119 d) .0359 C)1587 distribution of all possible samples of size 1. The mean of the sample means is equal to shan 1 standand deviation from the meas and asamping 295 e) .025
Answer #1

17)

According to the z score table

P(Z > 0.67) = 0.2514

18)

P(-1.25<Z < 1.0) = P(Z < 1)- P(Z < -1.25)

Option D 0.7357 is correct

19)

Z = (X - μ) / σ
Z = (88 - 100) / 15

Z = -0.8

According to the Z score table

P( X < 88) = 0.2119

20)

At greater than 1 standard deviation from mean Z = 1

According to Z score table

P( Z > 1) = 0.1587

Option D 0.1587 is correct

21)

Mean will be constant 2.95